OpenStudy (anonymous):

Is it possible to solve this? 3/7(2p-14)>6/7p-p

OpenStudy (anonymous):

yes it is

OpenStudy (anonymous):

How so?

OpenStudy (anonymous):

Fist you add p to both sides. Write the answer for me to to check.

OpenStudy (anonymous):

No luck? \[\frac{3}{7}(2p-14)+p>\frac{6}{7p} \] unless you made a mistake.

OpenStudy (anonymous):

Did you write the question right? Is it: \[\frac{3}{7}(2p-14)>\frac{6}{7p} -p\qquad?\] Or is it: \[\frac{3}{7}(2p-14)>\frac{6}{7}p -p\qquad?\]

OpenStudy (anonymous):

Its the second one.

OpenStudy (anonymous):

I just can't figure out how to do the first part.

OpenStudy (anonymous):

I multiply 3/7(2p), right?

OpenStudy (anonymous):

Yes and also \(\cfrac{3}{7}(-14)\)

OpenStudy (anonymous):

could you help me with that?